# The EBLM Project XII. An eccentric, long-period eclipsing binary with a companion near the hydrogen-burning limit

### v1

@@ Abstract
In the hunt for Earth-like exoplanets it is crucial to have reliable host star parameters, as they have a direct impact on the accuracy and precision of the inferred parameters for any discovered exoplanet. For stars with masses between 0.35 and 0.5 ${\rm M_{\sun}}$ an unexplained radius inflation is observed relative to typical stellar models. However, for fully convective objects with a mass below 0.35 ${\rm M_{\sun}}$ it is not known whether this radius inflation is present as there are fewer objects with accurate measurements in this regime. Low-mass eclipsing binaries present a unique opportunity to determine empirical masses and radii for these low-mass stars. Here we report on such a star, EBLM J2114-39 B. We have used HARPS and FEROS radial-velocities and TESS photometry to perform a joint fit of the data, and produce one of the most precise estimates of a very low mass star’s parameters. Using a precise and accurate radius for the primary star using Gaia DR3 data, we determine J2114-39 to be a $M_{1}=0.987\pm 0.059$ ${\rm M_{\sun}}$ primary star hosting a fully convective secondary with mass $M_{2}~{}=~{}0.0986~{}\pm 0.0038~{}\,\mathrm{M_{\sun}}$, which lies in a poorly populated region of parameter space. With a radius $R_{2}=~{}0.1275~{}\pm 0.0020~{}\,\mathrm{R_{\sun}}$, similar to TRAPPIST-1, we see no significant evidence of radius inflation in this system when compared to stellar evolution models. We speculate that stellar models in the regime where radius inflation is observed might be affected by how convective overshooting is treated.
@@ 2 Observations
3 Spectroscopic Analysis and Stellar Properties
Atmospheric parameters for the primary star were obtained using the PAWS pipeline as part of the work completed in Freckelton et al. (2023) which is based on iSpec (Blanco-Cuaresma, 2019). One strength of PAWS is that effective temperature ($T_{\rm eff}$), surface gravity ($\log\,g_{1}$), metallicity ($\rm[Fe/H]$), microturbulent velocity ($v_{\rm mic}$), and project rotational velocity ($vsini_{\star}$) are fitted for, when other commonly used methods will keep at least one fixed. Within the pipeline, initial estimates of $T_{\rm eff}$, $\log\,g_{1}$, $\rm[Fe/H]$, and $v_{\rm mic}$ were generated using the curve-of-growth equivalent widths method with the WIDTH (Sbordone et al., 2004) radiative transfer code. PAWS then uses these as inputs to a spectral synthesis method, wherein free parameters are iterated until a synthesised spectrum based upon them matches the observed one. The synthetic spectra are generated using the SPECTRUM radiative transfer code (Gray & Corbally, 1994). $T_{\rm eff}$, $\log\,g_{1}$, $\rm[Fe/H]$, and $v_{\rm mic}$ were all set as free parameters during the analysis of J2114-39. Employment of the spectral synthesis method means we are also able to include macroturbulent velocity ($v_{\rm mac}$) and projected rotational velocity ($v\sin i_{1}$) as free parameters.
As part of the pipeline all spectra are shifted into the laboratory frame, and normalised to a continuum flux of 1.0. In order to produce the best results possible, the 16 HARPS spectra for J2114-39 were co-added to achieve a signal-to-noise ratio, $SNR=257$. The pipeline assumes solar-type parameters as an input to the equivalent widths method, using those outlined by Blanco-Cuaresma (2019). To ensure homogeneity throughout analysis, the SPECTRUM (Gray & Corbally, 1994) line list was used in both methods present in the pipeline, in addition to the ATLAS (Kurucz, 2005) set of model atmospheres.
The primary’s mass and radius are derived by interpolation of MIST isochrones (Dotter, 2016; Choi et al., 2016) based on the atmospheric parameters $\rm T_{\rm eff}$, $\rm[Fe/H]$, together with the parallax and infrared colours as input parameters and using a nested sampling approach implemented in the isochrones package (Morton, 2015). The stated uncertainty is the average of the errors calculated from the 16 and 84 percentiles of the resulting distribution. All values, taken from Freckelton et al. (2023) are reported in Table 2.
@@ 5 Parameter derivation
To extract accurate parameters on the secondary star, we need accurate parameters on the primary. Thanks to Gaia DR3 (Gaia Collaboration et al., 2016) the stellar radius ($R_{1}$) is better estimated than the mass ($M_{1}$) that traditionally relies on applying stellar models. However, thanks to our lightcurve modelling we can constrain $M_{1}$ using the stellar density. From Kepler’s law, stellar density ($\rho_{\rm 1}$) is defined as
@@ 6 Results and Conclusion
After the global modelling and parameter derivation, we determine a mass function $f_{m}=0.0008127\pm 0.0000028~{}\,\mathrm{M_{\sun}}$, indicating a low-mass companion. Then we calculate the secondary’s surface gravity $\log g_{2}=5.222\pm 0.0135$, which is found in absolute terms (Southworth et al., 2007) from parameters fit with the MCMC. This value alone confirms the secondary is a dense star, near the bottom of the main-sequence. Assuming parameters for the primary star (as in Section 3; $R_{1}=1.279\pm 0.017~{}\,\rm R_{\odot}$ and Section 5: $M_{1}=0.987\pm 0.059~{}\,\rm M_{\odot}$), we find the secondary stellar companion to be a late M-dwarf with a mass $M_{2}~{}=~{}0.0986~{}\pm 0.0038~{}\,\mathrm{M_{\sun}}$and a radius $R_{2}=~{}0.1275~{}\pm 0.0020~{}\,\mathrm{R_{\sun}}$and thus a mass ratio $q=0.0999\pm 0.0022$ with the primary. If we recalculate the surface gravity of the primary star using the obtained mass and radius, we find that the value is consistent within 1 sigma with the spectroscopic surface gravity, regardless of the choice of primary mass value. From the fit, we also find the system has a clearly detected orbital eccentricity $e=0.56767\pm 0.00045$, which is not unusual for a binary with this period (e.g. Triaud et al., 2017).
Without following the parameter derivationprocedure described in Section 5, we would have calculated the mass of the secondary star through equation 4 where we would have assumed the denominator to be $M_{1}^{2}$ as commonly done when $M_{2}\ll M_{1}$. By using the primary mass value calculated from isochrones, we would then have found the mass of the secondary companion to be $M_{2}=0.1038~{}\pm 0.0067~{}\,\rm M_{\odot}$ instead of $M_{2}~{}=~{}0.0986~{}\pm 0.0038~{}\,\mathrm{M_{\sun}}$. Although these results are consistent within error margins (as can be seen in Fig. 3) this represents a $\sim 5.0\%$ systematic bias towards higher masses (and would therefore bias the overall population towards less inflated objects). Appendix A details why the fractional uncertainty on mass is smaller for the secondary star than for the primary.
@@
Table 2: Parameters of EBLM J2114-39. ${}^{(*)}$ from Freckelton et al. (2023) and ${}^{(**)}$ from GAIA DR3. Primary mass values given from the MIST isochrones $M_{1,\rm iso}$, as well as from the method described in Section 5. The quoted upper limits correspond to the 99.7 percentile of the posterior distribution.
@@
Primary $*{}^{*}$ $M_{1,\rm iso}$ ($\mathrm{M_{\odot}}$) $1.038\pm 0.064$
Primary mass $M_{1}$ ($\mathrm{M_{\odot}}$) $0.987_{-0.060}^{+0.058}$
Primary $*{}^{*}$ $R_{1}$ ($\rm R_{\odot}$) $1.279\pm 0.017$
Effective $*{}^{*}$ $T_{\rm eff}$ (K) $5763\pm 216$
$*{}^{*}$ [Fe/H] (dex) $-0.03\pm 0.14$
Spectroscopic surface $*{}^{*}$ $\log g_{1}$ (cgs) $4.241\pm 0.035$
Projected rotational $*{}^{*}$ $v\sin i_{1}$ ($\rm km\,s^{-1}$) $1.60\pm 0.22$
Macroturbulent $*{}^{*}$ $v_{\rm mac}$ ($\rm km\,s^{-1}$) $4.26\pm 1.37$
Microturbulent $*{}^{*}$ $v_{\rm mic}$ ($\rm km\,s^{-1}$) $1.08\pm 0.22$
${}^{**}$ Plx (mas) $4.30\pm 0.02$
@@
Secondary mass $M_{2}$ ($\mathrm{M_{\odot}}$) $0.0986_{-0.0039}^{+0.0037}$
Secondary radius $R_{2}$ ($\rm R_{\odot}$) $0.1275\pm 0.0020$
Mass Ratio $q$ $0.0999_{-0.0020}^{+0.0023}$
Impact parameter $b$ $<0.348$
Total eclipse duration $T_{\mathrm{tot}}$ (h) $11.624_{-0.097}^{+0.11}$
Full eclipse duration $T_{\mathrm{full}}$ (h) $9.469_{-0.097}^{+0.11}$
Eclipse depth $\delta_{\mathrm{undil}}$ $0.01140_{-0.00015}^{+0.00016}$
@@
Figure 3: Mass-radius diagram with logarithmic mass axis, showing published single-lined systems (SB1) in grey, double-lined binaries (SB2) in black and EBLM J2114-39 in coral colour. The $1~{}\rm Gyr$ and $5~{}\rm Gyr$ isochrones are shown in purple and teal respectively. The lower panel highlights differences between the objects and the $5~{}\rm Gyr$ isochrone. Black and grey data points from compendium (and cited papers) in Triaud et al. (2020), with additional points from Morales et al. (2009).
@@ Acknowledgements
This research is supported from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement ${}^{\circ}$ 803193/BEBOP), and by a Leverhulme Trust Research Project Grant (${}^{\circ}$ RPG-2018-418).
@@ Appendix A Error calculation for stellar masses
The fractional error on the secondary star is smaller than the fractional error on the primary mass. This may be surprising, but is actually a direct result of the fit to the RVs being very good and the different powers in the mass ratio of equation 4. From equation 5, we can write that
@@
The error in $M$ can be calculated from equation 6:

### v2

@@ Abstract
In the hunt for Earth-like exoplanets it is crucial to have reliable host star parameters, as they have a direct impact on the accuracy and precision of the inferred parameters for any discovered exoplanet. For stars with masses between 0.35 and 0.5 ${\rm M_{\sun}}$ an unexplained radius inflation is observed relative to typical stellar models. However, for fully convective objects with a mass below 0.35 ${\rm M_{\sun}}$ it is not known whether this radius inflation is present as there are fewer objects with accurate measurements in this regime. Low-mass eclipsing binaries present a unique opportunity to determine empirical masses and radii for these low-mass stars. Here we report on such a star, EBLM J2114-39 B. We have used HARPS and FEROS radial-velocities and TESS photometry to perform a joint fit of the data, and produce one of the most precise estimates of a very low mass star’s parameters. Using a precise and accurate radius for the primary star using Gaia DR3 data, we determine J2114-39 to be a $M_{1}=0.998\pm 0.052$ ${\rm M_{\sun}}$ primary star hosting a fully convective secondary with mass $M_{2}~{}=~{}0.0993~{}\pm 0.0033~{}\,\mathrm{M_{\sun}}$, which lies in a poorly populated region of parameter space. With a radius $R_{2}=~{}0.1250~{}\pm 0.0016~{}\,\mathrm{R_{\sun}}$, similar to TRAPPIST-1, we see no significant evidence of radius inflation in this system when compared to stellar evolution models. We speculate that stellar models in the regime where radius inflation is observed might be affected by how convective overshooting is treated.
@@ 2 Observations
3 Stellar Properties
We used the HARPS spectra to determine the stellar atmospheric parameters. The 16 available spectra for J2114-39 were shifted into the laboratory frame and normalised to a continuum level of 1.0. The spectra were then co-added to achieve a combined signal-to-noise ratio, $SNR=257$. The following atmospheric parameters for the primary star were derived using iSpec (Blanco-Cuaresma et al., 2014; Blanco-Cuaresma, 2019): the effective temperature, $T_{\rm eff}$, surface gravity, $\log\,g_{1}$, metallicity, $\rm[Fe/H]$, and microturbulent velocity, $v_{\rm mic}$. We applied the curve-of-growth equivalent widths method with both the WIDTH (Sbordone et al., 2004) and MOOG (Sneden et al., 2012) radiative transfer codes separately. Solar parameters were input, using those outlined by Blanco-Cuaresma (2019). The SPECTRUM (Gray & Corbally, 1994) line list was used in addition to the ATLAS (Kurucz, 2005) set of model atmospheres. To reduce systematic error, the final stellar atmospheric parameters displayed in Table 2 are weighted averages of the results from both the WIDTH and MOOG code.
The primary’s mass and radius are derived by interpolation of MIST isochrones (Dotter, 2016; Choi et al., 2016) based on the atmospheric parameters $\rm T_{\rm eff}$, $\rm[Fe/H]$, together with the parallax and infrared colours as input parameters and using a nested sampling approach implemented in the isochrones package (Morton, 2015). The stated uncertainty is the average of the errors calculated from the 16 and 84 percentiles of the resulting distribution. All values are reported in Table 2.
@@ 5 Parameter derivation
To extract accurate parameters on the secondary star, we need accurate parameters on the primary. Thanks to Gaia DR3 (Gaia Collaboration et al., 2016) the stellar radius ($R_{1}$) is better estimated than the mass ($M_{1}$) that traditionally relies on applying stellar models. However, with our lightcurve modelling we can constrain $M_{1}$ using the stellar density. From Kepler’s law, stellar density ($\rho_{\rm 1}$) is defined as
@@ 6 Results and Conclusion
After the global modelling and parameter derivation, we determine a mass function $f_{m}=0.0008127\pm 0.0000028~{}\,\mathrm{M_{\sun}}$, indicating a low-mass companion. Then we calculate the secondary’s surface gravity $\log g_{2}=5.222\pm 0.0135$, which is found in absolute terms (Southworth et al., 2007) from parameters fit with the MCMC. This value alone confirms the secondary is a dense star, near the bottom of the main-sequence. Assuming parameters for the primary star (as in Section 3; $R_{1}=1.283\pm 0.012~{}\,\rm R_{\odot}$ and Section 5: $M_{1}=0.998\pm 0.052~{}\,\rm M_{\odot}$), we find the secondary stellar companion to be a late M-dwarf with a mass $M_{2}~{}=~{}0.0993~{}\pm 0.0033~{}\,\mathrm{M_{\sun}}$  and a radius $R_{2}=~{}0.1250~{}\pm 0.0016~{}\,\mathrm{R_{\sun}}$  and thus a mass ratio $q=0.0995\pm 0.0019$  with the primary. If we recalculate the surface gravity of the primary star using the obtained mass and radius, we find that the value is consistent within 1 sigma with the spectroscopic surface gravity, regardless of the choice of primary mass value. From the fit, we also find the system has a clearly detected orbital eccentricity $e=0.56733\pm 0.00027$, which is not unusual for a binary with this period (e.g. Triaud et al., 2017).
Without following the parameter derivation procedure described in Section 5, we would have calculated the mass of the secondary star through equation 4 where we would have assumed the denominator to be $M_{1}^{2}$ as commonly done when $M_{2}\ll M_{1}$. By using the primary mass value calculated from isochrones, we would then have found the mass of the secondary companion to be $M_{2}=0.0999~{}\pm 0.0045~{}\,\rm M_{\odot}$ instead of $M_{2}~{}=~{}0.0993~{}\pm 0.0033~{}\,\mathrm{M_{\sun}}$. Although these results are consistent within error margins (as can be seen in Fig. 3) this represents a $\sim 1.0\%$ systematic bias towards higher masses (and would therefore bias the overall population towards less inflated objects). Appendix A details why the fractional uncertainty on mass is smaller for the secondary star than for the primary.
@@
Table 2: Parameters of EBLM J2114-39. (∗) from Sec. 3, and (∗∗) from GAIA DR3. Primary mass values given from the MIST isochrones $M_{1,\rm iso}$, as well as from the method described in Section 5. The quoted upper limits correspond to the 99.7 percentile of the posterior distribution.
@@
Primary mass $M_{1,\rm iso}$ ($\mathrm{M_{\odot}}$) $1.003\pm 0.042$
Primary mass $M_{1}$ ($\mathrm{M_{\odot}}$) $0.998_{-0.054}^{+0.049}$
Primary radius $R_{1}$ ($\rm R_{\odot}$) $1.283\pm 0.012$
Effective temperature $T_{\rm eff}$ (K) $5667\pm 113$
Metallicity [Fe/H] (dex) $-0.001\pm 0.032$
Spectroscopic surface gravity $\log g_{1}$ (cgs) $3.936\pm 0.010$
Microturbulent velocity $v_{\rm mic}$ ($\rm km\,s^{-1}$) $0.839\pm 0.010$
Parallax∗∗ Plx (mas) $4.30\pm 0.02$
@@
Secondary mass $M_{2}$ ($\mathrm{M_{\odot}}$) $0.0993_{-0.0035}^{+0.0031}$
Secondary radius $R_{2}$ ($\rm R_{\odot}$) $0.1250\pm 0.0016$
Mass Ratio $q$ $0.0995_{-0.0017}^{+0.0020}$
Impact parameter $b$ $<0.120$
Total eclipse duration $T_{\mathrm{tot}}$ (h) $11.624_{-0.097}^{+0.108}$
Full eclipse duration $T_{\mathrm{full}}$ (h) $9.469_{-0.097}^{+0.109}$
Eclipse depth $\delta_{\mathrm{undil}}$ 0.01141±0.000170.01141$\pm$-0.000170.01141 ± - 0.00017
@@
Figure 3: Mass-radius diagram with logarithmic mass axis, showing published single-lined systems (SB1) in grey, double-lined binaries (SB2) in black and EBLM J2114-39 results overlapping in turquoise and coral colour. The $1~{}\rm Gyr$ and $5~{}\rm Gyr$ isochrones are shown in purple and teal respectively. The lower panel highlights differences between the objects and the $5~{}\rm Gyr$ isochrone. Black and grey data points from compendium (and cited papers) in Triaud et al. (2020), with additional points from Morales et al. (2009).
@@ Acknowledgements
This research is supported from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement n 803193/BEBOP), and by a Leverhulme Trust Research Project Grant (n RPG-2018-418).
@@ Appendix A Error calculation for stellar masses
The fractional error on the secondary star is smaller than the fractional error on the primary mass. This may be surprising, but is actually a direct result of the fit to the RVs being very good and the different powers in the mass ratio of Eqn 4. From Eqn 5, we can write that
@@
The error in $M$ can be calculated from Eqn 6: